OpenStudy (anonymous):

Taylor Series Question! f(x) = e^(-x)+sinx -1; I'm trying to propose a reformation using taylor series to clear up cancelation near x = 0. Please help!

OpenStudy (anonymous):

I don't understand why my answer to finding the series should be f(x) = [1+x + x^2/2!....]; I get the rest of the summation, but don't get where I should be getting "1+"

OpenStudy (anonymous):

B/c I know "x" comes from ((-1)^(1) * (x)(1))/1!

OpenStudy (anonymous):

I'm not sure I understand the question... Do you want to find the Taylor series of \(f(x)=e^{-x}+\sin x-1\) ?

OpenStudy (anonymous):

Yes, and I know some of the general summation which make solving it easier, like e^(-x) = SUM ((-1)^(n) * (x)^(n))/n!

OpenStudy (anonymous):

But I'm not getting my answer right.

OpenStudy (anonymous):

First off, I do know that the function is terminal at n =5, and then it starts to repeat. So I know I'm going to have a remainder term at n = 5, so I'm solving the taylor series for only up to that point. And I also know I have to solve e^(-x) separately and add it to the taylor series part of sinx, and then -1, in a string (I hope this makes since). But I don't know what I 'm doing that makes my answer wrong.

OpenStudy (anonymous):

\[e^{x}=\sum_{n=0}^\infty\frac{x^n}{n!}\] So \[e^{-x}=\sum_{n=0}^\infty\frac{(-x)^n}{n!}\] Also, \[\sin x=\sum_{n=0}^\infty\frac{(-1)^nx^{2n+1}}{(2n+1)!}\] So you have \[f(x)=\left(1-x+\frac{x^2}{2}-\frac{x^3}{3!}+\cdots\right)+\left(x-\frac{x^3}{3!}+\frac{x^5}{5!}-\cdots\right)-1\] Like you mentioned, all the odd powers past 5 and including 1 and 5 will disappear. \[f(x)=\left(1+\frac{x^2}{2}-\frac{x^3}{3!}+\frac{x^4}{4!}+\frac{x^6}{6!}+\cdots\right)+\left(-\frac{x^3}{3!}\right)-1\] \[f(x)=\frac{x^2}{2}-\frac{2x^3}{3!}+\frac{x^4}{4!}+\frac{x^6}{6!}+\cdots\]

OpenStudy (anonymous):

I'm not seeing a compact version of this, except for the even powers...

OpenStudy (anonymous):

Thanks so much for your reply! So, at n = 0, I'm doing -x^(0)/(0!), correct? Why am I not getting 1?

OpenStudy (anonymous):

But \(\dfrac{(-x)^0}{0!}=\dfrac{1}{1}=1\)

OpenStudy (anonymous):

That's my confusion. What do I put into x?

OpenStudy (anonymous):

I know e^(-0) = 1. But I thought I had to use the sum? Sorry, this may be a dumb questions, but I haven't looked at calc 2 for about 3 years...

OpenStudy (anonymous):

It doesn't matter what \(x\) is. You only have to worry about the index \(n\). Notice how we keep the \(x\)'s in all the other terms?

OpenStudy (anonymous):

Correction: \(x\) can't be 0, since \(0^0\) is indeterminate, but like I said, we're only worrying about the \(n\).

OpenStudy (anonymous):

Oh, so we treat it as 1 * x? So I can compute -x^(0) as -1^(0)/0!...no, that doesn't give me the same answer

OpenStudy (anonymous):

What is the answer you should be getting?

OpenStudy (anonymous):

I should get [1-x+x^2/2!....Ox^5]...(I understand the sinx part, I think). The O I think stands for my error/remainder part.

OpenStudy (anonymous):

Instead of O(X^5), I wrote (-\xi)^(5)/5!, which I think means the same thing.

OpenStudy (anonymous):

Hmm, my understanding of big O notation is rusty... Wolfram gives an answer different from that one.

OpenStudy (anonymous):

That's okay. I just need understanding about the first term. For my e^(-x), my first term is (-x)^(0)/0! How do I know how to calculate that? Or for the sinx term:" -x/1!---I'm supposed to get x out of that one, but I have -x. I don't know know what I'm doing wrong.